# Self-Organising Memristive Networks as Physical Learning Systems

### v1

@@ 3.2 Mean field theory of conductance transitions
Critical phenomena in SOMNs are closely linked to self-organisation and feedback dynamics, which enable adaptive states under external driving. These behaviours often appear as scale–free event distributions—such as avalanches [16]—and are key to understanding emergent correlations. In SOMNs, global conductance can exhibit sharp transitions with voltage or component density, arising from the interplay between local dynamics and global connectivity. The mechanisms behind these critical states warrant further theoretical investigation. Related studies in neuromorphic nanowire networks [93, 94, 95], percolating NP networks [75, 54], and stochastic switching systems [76] provide experimental and computational evidence for avalanche dynamics and phase transitions. Theoretical insights into such transitions have also been drawn from analogies to random fuse models [121], dielectric breakdown [98], and tunneling-dominated percolation [102], which are relevant to understanding the observed deviations from classical universality classes in SOMNs [58, 101].
Mean-field models have been developed for the percolation transition [99], drawing on approaches from random fuse models [121]. While such mean-field approaches offer insight into general system behaviour, they fall short in accurately predicting critical exponents, which require more refined analysis [122] and consideration of effects like tunneling that are normally not consisdered in percolation models [102]. Crucially, to be effective, mean-field theory must incorporate the “liquid” nature of SOMNs—their dynamic, state-dependent connectivity, which is central to capturing their collective and emergent properties.
@@
Figure 3: Modeling emergent behaviour of memristive networks. (a) Experimental data on potentiation and relaxation dynamics related to short-term plasticity effects in NW networks under voltage stimulation (blue line), grid-graph modeling (black dashed line) and mean-field theory results (green dashed line) obtained by using methods outlined in sec.  3. (b) Results from grid-graph modeling allow to visualize spatiotemporal dynamics of the network, characterized by the progressive formation of a conductive pathway bridging stimulated electrodes that progressively fades way after the end of stimulation. The red intensity is proportional to the conductance of edges, while the blue intensity is proportional to the node voltage. Adapted from Ref.  [123] and  [109].
@@
Using the theoretical framework outlined in sec. 3.1, recent work has formulated a dynamical mean-field theory that attempts to characterize the voltage-induced dynamics of SOMNs by reducing equations of the type (4) to a single mean field dynamical equation. This is written in terms of an effective potential $\mathcal{V}_{\Delta v}(\langle g\rangle)$, which governs the evolution of the network’s average conductance state [120, 109], and allows determination of whether the interacting system is in the low or high conducting phase. These phases are determined by the state of the junctions, e.g. all on or all off. The resulting equation of motion is of the form:
@@
where $\Delta v$ is the applied voltage. As voltage increases, the shape of $\mathcal{V}$ changes, potentially introducing multiple minima corresponding to different conducting states. A mean field equation of the type (7) has been validated experimentally [109] and Fig. 3 shows a comparison between experimental data on short-term synaptic plasticity effects and theoretical response curves obtained through mean-field theory, as well as simulation results based on a grid–graph model that allows to directly visualize spatiotemporal dynamics of the system [123].
@@
Figure 4: Conductance transitions in memristive networks. a. Schematic illustration of conductance phase transitions in SOMNs, where the macroscopic order parameter (average conductance $G$) depends on control parameters such as applied voltage $V$, nanowire/nanoparticle (areal) density $\rho$, or average connectivity $d$. The shaded region represents the transition regime where emergent collective behaviour manifests. Depending on system parameters and underlying dynamics, this transition can be continuous (second-order), abrupt (first-order), or gradual, reflecting different types of dynamical phase behaviour. b. Numerical simulation showing the time evolution of internal junction conductance states (red lines) across the network following a voltage step near the critical point, when the effective potential develops multiple competing minima. The black line shows the ensemble average $\langle g\rangle$ across all junctions, while the dashed blue line corresponds to the prediction from mean-field theory (MFT). The divergence between simulation and MFT near the critical regime illustrates the breakdown of mean-field approximations due to the onset of strong feedback and topological reconfiguration. The inset shows the subcritical regime (single minimum in the effective potential), where junction conductances decay smoothly (laminar dynamics) and MFT matches simulations well. These results highlight the emergence of collective, nonlinear dynamics and the importance of including feedback and stochastic effects beyond mean-field approaches.
@@
Various methods have been used to obtain an equation of the form (7) from eqn. (4). At the core of these methods is how the matrix inverse $(I\pm\chi\Omega_{B}\mathcal{G})^{-1}$ enters into eqn. (4). This matrix inverse, once expanded, leads to non-linear terms between the variables $g_{i}$’s, and thus couples the junctions memory states across the network, and its structure gives rise to feedback and phase transitions. The first method is based on random matrix theory averaging of the inverse while keeping fixed row and column sums  [120, 128, 129]. Another method involves preserving certain graph-theoretic properties (e.g., number of cycles), leading to orthogonal transformations $\Omega_{B}^{\prime}=O\Omega_{B}O^{T}$ (with $O$ an orthogonal matrix), enabling exact averaging over transformations via Weingarten calculus [130, 131]. Alternatively, and more brutally in [109], the mean field inverse has been obtained by minimising the Frobenius norm between the exact and the mean field inverse. The resulting effective potential shape depends on the voltage, which implies that conductance transitions between the low and high conductances are controlled by a macroscopic theory. When there is only a single minimum, the dynamics of the junction is laminar and the MFT works well numerically. However, the devil is in the details. Near transition points, e.g. when the effective potential develops multiple minima, the system becomes susceptible to noise, which can be amplified and can induce transitions between these two states (cf. Fig. 4b), and a failure of the MFT. Depending on the shape of the effective potential, these transitions can be both of the first or second order (continuous or discontinuous) type, dependent on voltage, in addition to static geometrical effects due to density (cf. Fig. 4a). Further studies and new techniques [132] are necessary to understand the behaviour of SOMNs as complex physical systems. It is important to stress that mean field theory is not effective near dynamical phase transitions; as we can see in Fig. 4b, the match between the dashed blue (MFT) and solid black line (simulations) is poor near the threshold between the laminar and turbulent regime, while it is quite remarkable in the laminar one (inset, subcritical). Thus, alternative approaches (e.g. self-organised criticality [133]) are necessary to describe such critical points.
@@
In summary, theoretical models provide a foundation for understanding how SOMNs generate the collective nonlinear dynamics observed experimentally. From lumped-element to mean-field approaches, these frameworks reveal how memory, phase transitions, and adaptive responses emerge. These emergent dynamics represent physical information encoding and form the basis of a physical learning substrate, especially near transitions. The next section explicitly shows how these dynamics enable learning.
@@ 4 Physical learning with SOMNs
Figure 5: Physical learning paradigms for self-organising memristive networks. Arbitrary input data (left) is first pre-processed and encoded into input voltage signals (centre left) and delivered to a multi-terminal SOMN device (centre main), where conductance changes adapt to the input signal patterns. Dynamical reconfiguration of the SOMN, governed by internal transport dynamics and Kirchhoff constraints, physically compresses signal information to enable learning (output). Two representative learning paradigms are shown (right). In reservoir computing, the SOMN’s high-dimensional dynamical features are read out (red solid curves) and exploited in an external machine learning model; feedback to the input (red dashed curve) is also applied in some cases (e.g. recursive prediction). In associative and contrastive learning, feedback is applied based on output (green dotted lines). Both approaches enable flexible, substrate-intrinsic learning using only input-output access, without any direct addressability of memristive components.
@@
With an understanding of the mechanisms that underlie adaptive dynamics in SOMNs, we now show how these properties can be operationalised in physical learning systems. To provide some context, it is important to note that solving complex learning tasks requires nonlinear integration, as became apparent after the perceptron was introduced  [9, 134]. Nonlinear integration can be achieved either through multilayer networks [11] and/or recurrent dynamics [10].
@@
Neuromorphic approaches aim to realise such nonlinearities by mimicking biological synapses and/or neurons, either through emulation or simulation [17]. The emulation approach involves exploiting material and devices (e.g. memristors) that natively exhibit neuromorphic nonlinear dynamics [135]. The simulation approach instead typically involves algorithms that simulate point neuron spike models [136]. While both approaches can mimic synaptic plasticity, which is generally regarded as the key mechanism underlying the brain’s unique learning abilities [137], how this is implemented at the circuit device level differs considerably. For example, synaptic conductances in memristor crossbar devices are often used to physically instantiate matrix–vector operations to perform inference on globally-trained ANN models. On the other hand, spike-based neuromorphic circuit devices typically implement local learning algorithms (e.g. spike-timing dependent plasticity, STDP).
@@
Here, we highlight how physical learning can be achieved with SOMNs by exploiting their local and global brain-like nonlinear dynamics (including memory), without requiring direct addressability or control of their memristive circuit components. Two key learning paradigms are described below and illustrated in Fig. 5: physical reservoir computing and associative learning. These learning paradigms demonstrate how the dynamical states of SOMNs, as complex physical systems, can be exploited for learning either in an external output layer, in the case of reservoir computing, or within the network itself, in the case of associative learning.
@@ 4.1 Physical reservoir computing
Reservoir computing is a machine learning paradigm where a reservoir (a recurrent neural network) maps an input signal into a higher dimensional manifold that is then analysed by an output layer, which is the only part that needs to be trained [138], making this an efficient learning scheme (the weights inside the neural network are fixed and not trained.)
@@
In the physical or in materia implementation, a non-linear physical system is exploited as the reservoir [139, 140, 141, 142, 143]. In this context, SOMNs have been demonstrated to satisfy the key requirements of a physical reservoir [47, 106, 45, 144, 145, 31, 123, 146, 80, 74, 81, 147], endowing i) high dimensionality, with internal dynamics made accessible via multiterminal device configurations; ii) high–order nonlinearity, realised through atomic / ionic transport mechanisms and recurrent network structure; iii) fading memory property (to ensure that reservoir states rely only on recent–past inputs and are independent of distant–past inputs) [106], as discussed in Sec. 2.2.
To date, physical reservoir computing with SOMNs has been experimentally demonstrated in a wide range of computing tasks. One of the earliest demonstrations was waveform regression using NW networks [45]. For NW-based SOMNs, these also include pattern recognition [123, 48], audio classification [47, 49], sequence memory [48] and waveform regression [45]. For SOMNs based on NPs physical reservoir computing has demonstrated time series prediction, non-linear transformation, waveform discrimination, memory capacity and speech recognition [74, 146], as well as speaker and Braille recognition [147]. Time series prediction and speech recognition was also experimentally achieved on NW-based SOMN in [148, 149], and theoretically investigated in [106]. In addition, the experimental results have been complemented by simulations that have been used to test in silico a range of issues including the scalability of these systems, the architecture of the reservoir[150, 74, 103], the impact of the positioning [92] and the number of electrodes and the dimensionality of the reservoir[110]. A wide range of computing tasks have been simulated including pattern recognition  [123, 151, 144], time series prediction  [149, 123, 145, 152], speech recognition  [148], regression [93, 31, 153] and to learn chaotic dynamics  [154]. Furthermore two-dimensional tracking and autonomous swarm generation has also been simulated [81].
The internal states of SOMNs and their emergent spatiotemporal dynamics [116] are both measured via multiterminal electrical measurements  [155, 61, 156]. These measurements are typically performed by measuring output currents or floating voltages, or by constructing a conductance matrix where each element corresponds to the conductance between a pair of electrodes. Reservoir computing can also be implemented in a two-terminal memristive network by adopting a time multiplexing scheme, as demonstrated for both NP [74] and NW networks [30].
Notwithstanding these attributes, physical reservoir computing is significantly different from algorithmic reservoir computing in the following ways: (i) physical reservoirs are constrained by physics and in the case of SOMNs, Kirchhoff’s electrical circuit conservation laws; (ii) SOMNs produce collective memory dynamics by virtue of their memristive elements; and (iii) physical reservoirs are dynamic networks, with non–static connectivity [157] For these reasons, physical reservoir computers cannot be mathematically described in the same way as algorithmic reservoir computers, which apply a mathematical nonlinear activation function solely to nodes in a recurrent network with a random, static connectivity structure, with many hyperparameters requiring optimisation. However, dynamics typical of SOMNs can be leveraged to develop new neural network models based on the reservoir computing paradigm [158, 154].
Similar to biological systems where functionalities are inherently related to the topology of neuronal networks [159], higher reservoir computing performance can be achieved using SOMNs with scale-free topology [74]. Higher computational performance can also be achieved experimentally via parallel reservoir architectures, which increase the reservoir dimensionality and richness of its output dynamics [146]. Importantly, it has been shown through simulations that performance in RC tasks using NW networks contacted with a restricted number of electrodes is close to the upper bounds achievable when using information from every NW [110], a key consideration for the experimental realisation of these systems, where information from each NW is not practically accessible. Such simulations provide additional insight into experimental findings that a large number of electrode readouts may not be necessary to achieve high performance, at least on some RC tasks implemented on NW network devices [155, 48]. Simulations and experimental implementations of more complex RC tasks are needed to qualify these results further.
@@
Additionally, it has been shown experimentally that the spatiotemporal information processing capabilities in multiterminal systems allow unconventional implementation of RC by exploiting the same electrodes as both reservoir inputs and readouts, allowing a reduction of hardware complexity without limiting computing capabilities [151]. Additionally, while the output layer is typically trained in software,
a fully memristive architecture can be realised by coupling SOMNs acting as a reservoir with memristive crossbar arrays [123, 160]. Furthermore, it has been revealed that memory patterns embedded in dynamical features of NW networks enable online learning and recall of spatiotemporal sequence patterns, thus showing the possibility of online learning from rich spatiotemporal dynamics generated by the SOMN, in contrast to data-driven statistical machine learning techniques [48].
@@
Several open questions remain in physical reservoir computing with SOMNs, including the final energy consumption for real world tasks, and whether methods can be developed for more self-contained learning that does not rely on training an external layer [2]. Indeed, such an approach could help bridge the gap between physical reservoir computing and associative learning, which is discussed in the next section.

### v2

@@ 3.2 Mean field theory of conductance transitions
Critical phenomena in SOMNs occur in at least two related but distinct forms. First, avalanche criticality has been observed experimentally in both NW and NP networks, where the electrical response exhibits scale–free switching statistics and long-range spatiotemporal correlations [16, 108, 109, 87, 74, 84] (see Fig. 3). In NP systems in particular, these behaviours are closely connected to transport near the percolation threshold, so percolation-based descriptions provide a natural starting point for interpreting the emergence of scale–free dynamics. Second, SOMNs appear to exhibit memory-driven conductance transitions that are not purely geometrical, but instead arise from the interplay between memristive internal states, external driving, and Kirchhoff-constrained network feedback [137, 116, 108]. In this case, the transition refers to the emergence and coexistence of multiple macroscopic conductance states, and to switching or bifurcation between low- and high-conductance regimes as control parameters such as voltage, density, or connectivity are varied. Analogies to random fuse models [138], dielectric breakdown [112], and tunneling-dominated percolation [117] remain useful for framing these behaviours, but they do not fully capture the role of memory in shaping such dynamical transitions. From this perspective, avalanche criticality and percolation are established experimental motifs in SOMNs, whereas these memory-induced conductance transitions are also of theoretical interest in their own right, because they point to a class of non-equilibrium collective phenomena that likely require new analytical and computational methods to be understood fully.
Numerical studies further suggest that these conductance transitions are reflected in changes of the global conductance patterns of the network. To clarify whether such critical points exist, how many macroscopic states can coexist, and what determines the order and universality of the transition, analytical mean-field approaches are particularly valuable. For example, mean-field descriptions of the percolation transition in memristive networks have been developed by approximating each junction as a threshold-based conductance switch [113], building on ideas from random fuse models [138, 117]. While such approaches capture useful aspects of the collective behaviour, they are not sufficient to predict critical exponents in SOMNs quantitatively, since this requires a more refined treatment of junction-level dynamics [139], of transport mechanisms such as tunneling, and more generally of the feedback between memory and Kirchhoff-constrained network dynamics. For this reason, these transitions are not only relevant for interpreting experiments, but also constitute a theoretical problem per se, motivating the development of novel methods capable of treating memory, disorder, and non-equilibrium feedback on equal footing.
@@
Figure 4: modelling emergent behaviour of memristive networks. (a) Experimental conductance measurements (blue) showing potentiation and relaxation dynamics related to short-term plasticity effects in NW networks under voltage pulse stimulation (red), corresponding grid-graph modelling (black dashed line) and mean-field theory results (green dashed line) obtained by using methods outlined in sec. 3. (b) Corresponding grid-graph modelling visualisation of spatiotemporal network dynamics, characterised by the progressive formation of a conduction path bridging electrodes that progressively fades way after the end of stimulation. Red intensity is proportional to edge conductance, while blue intensity is proportional to node voltage. Adapted from Ref. [140] and [124].
@@
Using the theoretical framework outlined in sec. 3.1, a dynamical MFT has been formulated that attempts to characterise the voltage-induced dynamics of SOMNs by reducing equations of the type (4) to a single mean-field dynamical equation. This is written in terms of an effective potential $\mathcal{V}_{\Delta v}(\langle g\rangle)$, which governs the evolution of the network’s average conductance state [137, 124], and allows determination of whether the interacting system is in the low or high conducting phase. These phases are determined by the state of the junctions, e.g. all on or all off. The resulting equation of motion is of the form:
@@
where $\Delta v$ is the voltage drop on the device. As applied voltage increases, the shape of $\mathcal{V}$ changes, potentially introducing multiple minima corresponding to different macroscopic conducting states corresponding to low- and high- conductance. A mean-field equation of the type (7) has been validated experimentally [124] and Fig. 4 shows a comparison between experimental data on short-term synaptic plasticity effects and theoretical response curves obtained through MFT, as well as simulation results based on a grid–graph model that allows direct visualization of the spatiotemporal dynamics [140].
@@
Figure 5: Conductance transitions in memristive networks. a. Schematic illustration of conductance phase transitions in SOMNs, where the macroscopic order parameter (average conductance $G$) depends on control parameters such as applied voltage $V$, nanowire/nanoparticle (areal) density $\rho$, or average connectivity $d$. The shaded region represents the transition regime where emergent collective behaviour manifests. Depending on system parameters and underlying dynamics, this transition may be continuous (second-order) or abrupt (first-order), reflecting different types of dynamical behaviours. b. Numerical simulation results showing the time evolution of internal junction conductance states (red lines) across a memristive network following a voltage step near the critical point, when the effective potential develops multiple competing minima. The black line shows the ensemble average $\langle g\rangle$ across all junctions, while the dashed blue line corresponds to the prediction from mean-field theory (MFT). The divergence between simulation and MFT near the critical regime illustrates the breakdown of mean-field approximations due to the onset of strong feedback and network reconfiguration. The inset shows the subcritical regime (single minimum in the effective potential), where junction conductances decay smoothly (laminar dynamics) and MFT matches simulations well. These results highlight the emergence of collective, nonlinear dynamics and the importance of including feedback and stochastic effects beyond mean-field approaches. Adapted from Ref. [137].
Various methods have been used to obtain an equation of the form (7) from eqn. (4). At the core of these approximation methods is how the matrix inverses of the form $(I\pm\chi\Omega_{B}\mathcal{G})^{-1}$, representing loosely the nonlinearity induced by Kirchhoff’s laws, enters into eqn. (4). This matrix inverse, once expanded, leads to nonlinear terms between the $g_{i}$ variables and thus couples the junctions memory states across the network, and its structure gives rise to feedback and dynamical phase transitions in the order parameter $G$. Depending on the shape of the effective potential, these transitions can be both of the first or second order (continuous or discontinuous) type, dependent on voltage, in addition to static geometrical effects due to density. This is depicted schematically in Fig. 5(a). Beyond continuous and discontinuous phase transitions, SOMNs can in principle exhibit more complex transition phenomenology (e.g. higher-order, discontinuous at higher derivatives), depending on microscopic dynamics and network heterogeneity; these are not shown in Fig. 5(a) for clarity.
@@
There are various existing theoretical methods to derive a mean field theory. The simplest method for obtaining the mean-field matrix inverse is based on a random matrix theory averaging of the inverse while keeping fixed row and column sums [137, 145, 146], corresponding to global information such as the total number of junctions. A more involved method allows preserving certain graph-theoretic properties (e.g., number of cycles), and using averages over random orthogonal transformations $\Omega_{B}^{\prime}=O\Omega_{B}O^{T}$ (with $O$ an orthogonal matrix), enabling exact averaging over transformations via Weingarten calculus [147, 148]. Alternatively, and more brutally in [124], the mean-field inverse has been obtained by minimising the Frobenius norm between the exact and the mean-field inverse. The resulting effective potential shape depends on the voltage, which implies that conductance transitions between the low and high conductances are controlled by a macroscopic theory. When there is only a single minimum, the dynamics of the junction is laminar (a relaxational dynamics with negative Lyapunov exponents) and the MFT works well numerically. However, near transition points, e.g. when the effective potential develops multiple minima, the system becomes susceptible to noise, which can be amplified and can induce transitions between these two states and a failure of the MFT, as shown in Fig. 5(b).
In the transition regime, the Lyapunov exponent associated with the state trajectories becomes positive, implying a turbulent or possibly transiently chaotic regime [149]. As MFT is not effective near dynamical transitions, alternative approaches (e.g. self-organised criticality [150]) and new techniques [151] are needed to better understand the emergent behaviour of SOMNs near these critical regimes.
@@
Thus, theoretical models provide a foundation for understanding how SOMNs generate the collective nonlinear dynamics observed experimentally. From lumped-element to mean-field approaches, these frameworks reveal how memory, phase transitions, and adaptive responses emerge. These emergent dynamics represent physical information encoding and form the basis of a physical learning substrate, especially near transitions. The next section explicitly shows how these dynamics enable learning.
@@ 4 Physical learning with SOMNs
Figure 6: Physical learning paradigms for self-organising memristive networks. Arbitrary input data (left) is converted into voltage signals and delivered to a multi-terminal SOMN device (centre main), where local conductance states of nanoscale memristive junctions adapt to the input signal patterns. Dynamic reconfiguration of the SOMN, governed by internal transport dynamics (ion migration, atomic filaments/hillocks, electron tunnelling) and constrained by Kirchhoff’s laws in a recurrent network circuitry, performs physical nonlinear transformation of signals to enable learning (output). Two representative learning paradigms are shown (right): in reservoir computing, the SOMN’s high-dimensional dynamical features are read out (green solid curves) and exploited in an external machine learning model (weights W); in associative learning, feedback is applied locally based on output (green dotted lines); this can be either at the input or the output or both, via applied voltage changes and electrode selection. Both approaches enable physical substrate-based learning using only input-output access, without any direct addressability of memristive components or global training schemes.
@@
With an understanding of the physical mechanisms that underlie adaptive dynamics in SOMNs, we now show how these properties can be exploited in physical learning systems. To place this in context, it is useful to recall that solving nontrivial learning tasks requires more than simple linear transformations, as became clear after the introduction of the perceptron [9, 152]. In practice, this expressive power can be achieved through multilayer architectures [11], recurrent dynamics [10], or combinations of both.
@@
Neuromorphic approaches aim to realise such nonlinearities by mimicking biological synapses and/or neurons, either through emulation or simulation [21]. The emulation approach involves exploiting materials and devices (e.g. memristors) that natively exhibit brain-like nonlinear dynamics [153]. The simulation approach instead typically involves algorithms that simulate point neuron spike models [154]. While both approaches can mimic synaptic plasticity, which is generally regarded as the key mechanism underlying the brain’s unique learning abilities [155], how this is implemented at the circuit and device level differs considerably. For example, synaptic conductances in memristor crossbar devices are often used to physically instantiate matrix–vector operations to perform inference on globally-trained ANN models. On the other hand, spike-based local learning algorithms (e.g. spike-timing dependent plasticity, STDP) have been implemented in spiking neural network (SNN) models [156] and neuromorphic hardware [157, 158].
@@
Here, we highlight how physical learning can be achieved with SOMNs by exploiting their local and global brain-like nonlinear dynamics (including memory), without requiring direct addressability or control of their circuit components. Furthermore, as SOMNs embed networks into their physical hardware, complex neural network algorithms are not required for learning (i.e. “the hardware is the software” [159]). Two key learning paradigms are described below and illustrated in Fig. 6: physical reservoir computing and associative learning. These learning paradigms demonstrate how the dynamical states of SOMNs, as complex physical systems, can be exploited for learning either in an external output layer, in the case of reservoir computing, or within the network itself, in the case of associative learning.
@@ 4.1 Physical reservoir computing
Reservoir computing is a machine learning paradigm where a reservoir (a recurrent neural network) maps an input signal into a higher dimensional manifold that is then analysed by a single output layer, which is the only part that needs to be trained [160], making this an efficient learning scheme (the weights inside the recurrent neural network are fixed and not trained, while a nonlinear activation function is applied to the nodes). Thus, the reservoir may be thought of as extracting features from the data which are then used to learn a model (e.g. regression, classification) in the output layer; the trained model is then used to perform inference on new features generated by the reservoir when presented with unseen data.
@@
In the physical or in materia implementation, a nonlinear physical system is used as the reservoir [161, 162, 163, 164, 165]. In this context, SOMNs have been demonstrated to satisfy the key requirements of a physical reservoir [54, 121, 52, 97, 166, 33, 140, 167, 90, 86, 91, 168], endowing i) high dimensionality, with internal dynamics made accessible via multiterminal device configurations; ii) high–order nonlinearity, realised through atomic / ionic transport mechanisms and recurrent network structure; iii) fading memory property (ensuring reservoir states rely only on recent–past inputs and are independent of distant–past inputs) [121], as discussed in Sec. 2.2. These properties ensure physical reservoirs can generate rich dynamical features from input signals with sufficient expressivity to fit a model to be learned in the output layer.
In SOMNs, fading memory arises naturally as a consequence of short-term memory of these systems [99, 140, 169, 167].
The internal states of SOMNs and their emergent spatiotemporal dynamics [132] are measured via multiterminal electrical measurements [170, 70, 61]. These measurements are typically performed by measuring output currents or floating voltages, or by constructing a conductance matrix where each element corresponds to the conductance between a pair of electrodes. Fig. 7 shows an example of the input–output mapping by a SOMN reservoir device. In this case, the inputs are sample images from the MNIST handwritten digit dataset that are converted into temporal voltage pulses (red). The output voltages (purple) are read out from device channels (channels 2 and 15 in this example) which produce expressive features for each digit image, with sufficient inter–digit separation to enable classification; the conductance output (green), obtained from the output current, encodes memory within the system and thus can also be leveraged in sequency memory tasks [55].
@@
Additionally, it has been shown that the spatiotemporal information processing capabilities in multiterminal devices allow an unconventional implementation of physical reservoir computing by exploiting the same electrodes as both inputs and readouts, thereby reducing hardware complexity without limiting computing capabilities [171]. Furthermore, while the output layer is typically trained in software, a fully memristive architecture can be realised by coupling SOMNs acting as a reservoir with memristive crossbar arrays [140, 172]. Physical reservoir computing can also be implemented in a two-terminal device by adopting a time multiplexing scheme [86, 32].
To date, physical reservoir computing with SOMNs has been experimentally demonstrated across a wide range of tasks, but often exploiting somewhat different operating regimes and readout modalities. Tasks that have demonstrated include waveform regression [52], image classification [140, 55], audio classification [54, 173, 56], time series prediction [86, 167, 174], non-linear transformations and memory capacity [163], and sequence memory [55], as well as speaker and Braille classification [168].
In addition, these experimental results have been complemented by simulations that have been used to test in silico various aspects including scalability, the reservoir architecture [175, 86, 101], the impact of electrode positioning [176], and the number of electrodes and the dimensionality of the reservoir[125]. A wide range of computing tasks have been simulated including pattern recognition [140, 171, 97], time series prediction [174, 140, 166, 177], speech recognition [173], regression [108, 33, 178] and prediction of multidimensional chaotic time series [179]. Furthermore, two-dimensional tracking and autonomous swarm generation has also been simulated [91].
Figure 7: Input–output feature mapping for MNIST classification and sequence memory with a nanowire-network physical reservoir. Column 1 shows average images of the digits 0, 3, 5, and 9, each obtained from 100 randomly selected MNIST training samples. Column 2 (red) shows the corresponding temporally encoded input-voltage pulse streams. Columns 3 and 4 (purple) show example voltage readouts from two device channels, which provide digit-dependent dynamical features for the classifier. Column 5 (green) shows the effective conductance $G$, which reflects the evolving internal memory state of the network and is useful for memory-related tasks, rather than classification itself. The figure therefore illustrates how the reservoir transforms input digits into discriminative dynamical features. Figure reproduced from Ref. [55].
Notwithstanding these attributes, physical reservoir computing with SOMNs differs significantly from algorithmic reservoir computing in the following ways: (i) SOMNs produce collective memory dynamics by virtue of their memristive elements; and (ii) SOMN reservoirs are dynamic networks, with non–static connectivity [180]; (iii) SOMN reservoirs are constrained by Kirchhoff’s electrical circuit conservation laws. For these reasons, SOMN reservoirs cannot be described mathematically in the same way as standard algorithmic reservoir computers. In conventional reservoir computing, nonlinearity is typically introduced through prescribed activation functions acting on the nodes of a recurrent network with usually random, static connectivity, and performance depends on the optimisation of multiple hyperparameters. By contrast, SOMN reservoirs are governed not only by memristive nonlinearities, but also by Kirchhoff’s circuit laws and, more generally, by the physical constraints imposed by the underlying material network and device geometry. Their dynamics therefore emerge from the coupled evolution of internal memory states, transport processes, and constrained network feedback, rather than from an abstract node-based update rule alone. In this sense, SOMNs are not simply hardware implementations of existing reservoir models, but physical systems whose native dynamics can inspire new neural-network models and extensions of the reservoir-computing paradigm [181, 179].
Similar to biological systems where functionalities are inherently related to the topology of neuronal networks [182], higher reservoir computing performance may be achieved using SOMNs with scale-free topology [86]. Higher computational performance can also be achieved experimentally via parallel reservoir architectures, which increase the reservoir dimensionality and richness of its output dynamics [167]. Importantly, it has been shown through simulations that performance in reservoir computing tasks using SOMNs contacted with a restricted number of electrodes is close to the upper bounds achievable when using information from every node [125], a key consideration for the experimental realisation of these systems, where information from each node is not practically accessible. Such simulations provide additional insight into experimental findings that a large number of electrode readouts may not be necessary to achieve high performance, at least on some tasks [170, 55, 169]. Simulations and experimental implementations of more complex reservoir computing tasks are needed to qualify these results further.
@@
Several open questions remain in physical reservoir computing with SOMNs, including the energy consumption for real world tasks, and whether methods can be developed for more self-contained learning that does not rely on training an external layer [2]. Indeed, such an approach could help bridging the gap between physical reservoir computing and associative learning, which is discussed in the next section.